If 60 ml of 0.3 F HAc is mixed with 200 ml of 0.1 F KOH. What is the hydronium ion concentration of the resulting solution?What is the pH of the solution?



1
Expert's answer
2021-03-22T06:12:43-0400

HAc + KOH = KAc + H2O

n (HAc) = n (KOH)

CM = n/V

n = CM x V

n (HAc) = 0.3 x 0.06 = 0.018 mol

n (KOH) = 0.1 x 0.2 = 0.02 mol

In the current conditions n (KOH) is greater than n (HAc). Therefore, all H ion is consumed and pH=0.